A cynlindrical water tank has a volume of 6(pie)x(to the power of 2)y(to the power of 4) cubic meters. The formula for the volume of a cylinder is (pie)r(to the power of 2)h. The water tank has a radius of xy meters. What is its height?
Answer by nerdybill(7384)

6(pi)x^2y^4 = (pi)r^2h
r was given as xy
.
6(pi)x^2y^4 = (pi)(xy)^2h
Solving for h:
6(pi)x^2y^4 = (pi)(xy)^2h
6x^2y^4 = (xy)^2h
6x^2y^4 = x^2y^2h
6y^4 = y^2h
6y^2 = h